OpenStudy (anonymous):

Medals! decide whether the equation is a trigonometric identity 1.sin^2 theta csc^2 theta=sin^2=cos^2 theta

OpenStudy (jdoe0001):

\(\large \bf sin^2(\theta)csc^2(\theta)=sin^2(\theta)cos^2(\theta)?\)

OpenStudy (anonymous):

csc(x) = 1/sin(x) \[\large \sin^{2}x (\frac{ 1 }{ \sin^{2}x }) = \sin^{2} x = \cos^{2} x\] \[\large 1 = \sin^2x - \cos^2 x? \] nope because \[\large \sin^2x + \cos^2 x =\] is an identity so the above statement is not true.

OpenStudy (anonymous):

okay, im following!

OpenStudy (anonymous):

\[\sin^2x + \cos^2 x = 1 \] correction ^^****

OpenStudy (anonymous):

so yeah, not a trig identity

OpenStudy (anonymous):

okay, thank you so much!